### v1

# PoGO: A Scalable Proof of Useful Work via Quantized Gradient Descent and Merkle Proofs

Abstract
We present a design called Proof of Gradient Optimization (PoGO) for blockchain consensus, where miners produce verifiable evidence of training large-scale machine-learning models. Building on previous work [1, 2, 3], we incorporate quantized gradients (4-bit precision [7] [8][9]) to reduce storage and computation requirements, while still preserving the ability of verifiers to check that real progress has been made on lowering the model’s loss. Additionally, we employ Merkle proofs over the full 32-bit model to handle large parameter sets and to enable random leaf checks with minimal on-chain data. We illustrate these ideas using GPT-3 (175B parameters) [5] as a reference example and also refer to smaller but high-performance models (e.g., Gemma 3 with 27B parameters). We provide an empirical cost analysis showing that verification is significantly cheaper than training, thanks in part to quantization and sampling. We also discuss the necessity of longer block times (potentially hours) when incorporating meaningful training steps, the trade-offs when using specialized GPU hardware, and how binary diffs may incrementally optimize updates. Finally, we note that fine-tuning can be handled in a similar manner, merely changing the dataset and the manner of sampling but preserving the overall verification flow. Our protocol allows verifiers to issue either positive or negative attestations; these are aggregated at finalization to either confirm the update or slash the miner.
1 Introduction
Traditional Proof-of-Work (PoW) blockchains (originating from Bitcoin [6]) rely on miners solving cryptographic puzzles that consume large amounts of energy without producing externally useful by-products. Recent Proof of Useful Work (PoUW) proposals [1, 2] have explored turning this computational effort into tasks beneficial to society, such as protein folding or machine-learning (ML) training. However, verifying that genuine ML computations have been performed (rather than possibly forged) poses a significant challenge due to the high dimensionality of modern models and the offline nature of large-scale training.
Proof of Gradient Optimization (PoGO) tackles this verification issue by requiring miners to commit to a new set of model parameters, $\mathbf{\theta}\in\mathbb{R}^{d}$, and prove that the update reduces the model’s loss
$\mathcal{L}(\mathbf{\theta})\;=\;\mathbb{E}_{(x,y)\in\mathcal{D}}\bigl{[}\ell\bigl{(}f_{\mathbf{\theta}}(x),\,y\bigr{)}\bigr{]},$
on a specified dataset $\mathcal{D}$. In practice, miners do a standard gradient-descent step
$\mathbf{\theta}_{t+1}\;=\;\mathbf{\theta}_{t}\;-\;\eta\,\nabla_{\mathbf{\theta}}\mathcal{L}(\mathbf{\theta}_{t}),$
for learning rate $\eta$. The protocol relies on:
  1. 1.
    Two-phase random verification of the quantized gradient and sample weights from the full model.
  2. 2.
    Quantized model publication (e.g., 4-bit weights) for lower-cost off-chain storage and verification [7, 8].
  3. 3.
    Merkle-based partial checks of the full 32-bit model so that large parameter sets (potentially gigabytes) need not reside entirely on-chain.
Together, these steps enable verifiers to efficiently check training correctness at scale, even for massive models like GPT-3 or Gemma 3.
Key contributions.
Our approach introduces several key innovations. First, we propose quantized gradient publication, where a 4-bit version of the model or updates is stored and transmitted off-chain. This reduces memory usage by roughly 8×8$\times$8 × while still preserving a 32-bit Merkle tree commitment for deeper verification. To ensure consistency between the quantized and full-precision models, we use Merkle-based random leaf verification: after a publicly verifiable random delay, the network requests a single randomly chosen leaf from the full-precision model’s Merkle tree to perform an efficient integrity check.
We also include an empirical cost analysis, estimating parameter counts and storage requirements for large models like GPT-3 (175B parameters) [5] as well as smaller, high-performance models such as Gemma 3 (27B parameters). Our findings show that verification is significantly cheaper than training, supporting a secure incentive model.
To address deployment realities, we examine block time and hardware trade-offs. Since training large models can take hours, block times may need to be longer. We analyze the cost dynamics between miners and verifiers, and show how 4-bit quantization and specialized hardware can offer up to an 8×8$\times$8 × speedup [7, 8].
We introduce an attestation mechanism where verifiers issue positive attestations when checks pass and negative ones when discrepancies are found [4]. These votes are collected in a final aggregator block that either finalizes the update or penalizes the miner. Finally, the protocol is compatible with fine-tuning tasks, which follow the same structure as full training with only minor adjustments.
2 Background and Motivation
2.1 Large-Scale ML Models on Blockchain
Modern large language models (LLMs) can have upwards of hundreds of billions of parameters. OpenAI’s GPT-3 has $\sim$175B parameters [5], which in 32-bit floating point can exceed 700 GB of raw parameter data. Meanwhile, smaller but still high-performance models (e.g., Gemma 3) might have around 27B parameters, which is significantly more compact (roughly 108 GB in 32-bit).
Storing or verifying such large models directly on-chain is infeasible. PoGO addresses this by using off-chain storage (e.g., IPFS) and on-chain commitments (hashes, Merkle roots), along with random sampling to ensure correctness. Furthermore, quantized versions of the weights (e.g., 4-bit) can reduce storage by a factor of 8, making distribution more practical [7, 8, 9].
2.2 Why Quantization?
Quantized models replace full-precision (e.g., 32-bit float) weights with lower-precision representations such as 4-bit integers. This significantly reduces:
  • Memory Footprint: 4×4$\times$4 × fewer bits than 16-bit, or 8×8$\times$8 × fewer than 32-bit.
  • Bandwidth Requirements: Cheaper to transmit off-chain or store in decentralized storage.
  • Compute Overheads: Specialized hardware (e.g., GPU tensor cores) can often process low-precision vectors at higher throughput, up to 8×8$\times$8 × faster for 4-bit vs. 32-bit [7, 9].
Hence, if PoGO requires public availability of model parameters, it makes sense to use a compact 4-bit representation. However, to preserve full precision for actual training and gradient checks, we still keep a Merkle commitment on the 32-bit model.
3 Protocol Overview
We outline PoGO’s core design in detailed form for each block at height $N$. Let $\theta_{t}\in\mathbb{R}^{d}$ represent the model parameters at iteration $t$, and suppose each block corresponds (conceptually) to an incremental update of the model using a gradient-based procedure. The timeline is illustrated in Figure 1, and each step is elaborated in subsequent sections.
  1. 1.
    Block N: Training and Commitment
    • A random miner (via VRF and based on stake) is selected to train a randomly chosen model from the list of active tasks.
    • The miner performs training steps (e.g., gradient descent) until it lowers the model loss $\mathcal{L}(\theta)$ in full precision (32-bit) more than a decrement $\epsilon$ predefined by the model owner. Formally, it must show
      $\mathcal{L}(\mathbf{\theta}_{t+1})\;<\;\mathcal{L}(\mathbf{\theta}_{t})-\epsilon.$
    • The miner also checks that the quantized (4-bit) version of the updated weights, call it $\widetilde{\theta}_{t+1}$, exhibits a lower loss on a small verification dataset (drawn from a VRF seed):
      ^$\widehat{\mathcal{L}}(\widetilde{\theta}_{t+1})\;<\;\widehat{\mathcal{L}}(\widetilde{\theta}_{t})\quad\text{on the random verification samples.}$.
[…]
    • The block includes:
      1. (a)
        hashFullModel32: The 32-bit model’s root Merkle hash.
      2. (b)
        hashQuant4: The 4-bit quantized model’s hash.
      3. (c)
        vrfProof: Proving the random choice of the training data mini-batch for verifying loss reduction.
  2. 2.
    Between Block N and N + (w/2)
    • The miner publishes the quantized (4-bit) model to IPFS or similar for data availability. This must happen by block $N+(w/2)$, where $w$ is the finalization window (e.g., 20 blocks).
    • Verifiers download and verify that hashQuant4 matches the published 4-bit model.
    • Verifiers also confirm that on the small verification dataset (seeded by the VRF), the quantized model indeed has lower loss than the previous model.
  3. 3.
    Block N + (w/2): Random Leaf Challenge & Merkle Proof
    • A new seed is derived from block $(N+w/2)$, which the miner cannot predict at block $N$.
    • From this seed, a random leaf of the 32-bit model’s Merkle tree is selected, say index $i$.
    • The miner must provide the corresponding leaf data $\texttt{leaf}_{i}$ and a Merkle proof linking it to hashFullModel32.
    • Verifiers check the leaf contents for consistency with the previously disclosed 4-bit quantized version (up to rounding). Any mismatch implies a faked or partially inaccurate full model.
  4. 4.
    After Block N + (w/2)
    • Each verifier issues either a positive attestation (if checks succeed) or a negative attestation (if a mismatch or failure is found). These are gossiped off-chain and then included in the final aggregator block at $N+w$.
    • The miner may optionally publish the entire 32-bit model in IPFS (or a portion, if using a sharding scheme). In some designs, only a fraction of verifiers need the full model; others might skip it if they trust the finalization.
  5. 5.
    Block N + w: Finalization and Possible Slashing
    • A new block leader aggregates all attestations from verifiers (positive or negative).
    • If $\geq 2/3$ of stake has signed with a positive attestation, the block is finalized (the update is accepted).
    • If the attestation threshold is not reached, or there are substantial negative attestations, the update is not accepted, and the miner is slashed a fraction of its stake for failing to produce a valid training step.
Block N:- Miner trains + - Publishes hashFullModel32, hashQuant4, vrfProofBetween N and N+(w/2):- Publish 4-bit model in IPFS- Verifiers confirm hashQuant4- Check quantized model’s lossBlock N+(w/2):- New random seed- Miner reveals random Merkle leaf- Verifiers check Merkle proofAfter N+(w/2):- Verifiers gossip (positive/negative) attestations- (Optional) 32-bit model publishBlock N+w:- Aggregator tallies attestations- $\geq 2/3$ positive $\implies$ update final- Otherwise, miner is slashed
Figure 1: Timeline of PoGO v2. The miner commits to the new full model and its quantized version at block N, publishes the 4-bit model by block $N+w/2$, and must reveal a random Merkle leaf (32-bit data) at block $N+w/2$. Verifiers issue positive or negative attestations, which are aggregated at block $N+w$ to finalize or slash.
4 Data Structures and Merkle Proofs
4.1 Merkle Tree for the 32-bit Model
A typical LLM can have billions of 32-bit parameters, easily reaching gigabytes in size. We chunk these parameters into leaves of size leafSize (e.g., 10MB). The final Merkle tree might have thousands or tens-of-thousands of leaves. The root of this tree, hashFullModel32, is published on-chain in block $N$.
4.2 Random Leaf Reveal
At block $(N+w/2)$, the protocol draws a public random seed from the chain state (e.g., a VRF from block $(N+w/2)$ itself). This seed is used to choose a random leaf index $i$. The miner reveals:
[…]
Each verifier checks:
  1. 1.
    The $\texttt{leaf}_{i}$ data matches the 4-bit quantized version in the corresponding region (allowing for rounding).
  2. 2.
    The Merkle path is valid.
A single leaf check is unlikely to catch a sophisticated forgery unless the miner tries to cheat across many parameters. However, PoGO’s two-phase sampling, plus the earlier proof that the quantized model actually lowers the loss on a random mini-batch, strongly reduces any cheating probability. If a verifier finds an inconsistency, it issues a negative attestation. To further strengthen security, the protocol could require multiple random leaves, each with a separate Merkle proof.
5 Quantized vs. Full-Precision Loss Checking
5.1 Training in 32-bit, Verifying in 4-bit
Miners conduct actual training (forward/backward passes) in full precision (32-bit) to preserve numerical stability. Then they produce a 4-bit version of the updated weights. The protocol requires that this 4-bit model also shows a lower loss on a small verification dataset. This ensures that the progress is “real” even in quantized form:
^$\widehat{\mathcal{L}}(\widetilde{\theta}_{t+1})\;<\;\widehat{\mathcal{L}}(\widetilde{\theta}_{t})\;-\;\epsilon_{\text{quant}},$
with $\epsilon_{\text{quant}}$ possibly slightly smaller than $\epsilon$ to account for quantization error.
5.2 Why This Works
Since 4-bit quantization can introduce rounding noise, it is possible that small improvements in 32-bit might disappear when converted to 4-bit. However, in practice, a well-chosen quantization method preserves enough accuracy to detect genuine loss improvements, especially if the improvement is not infinitesimal [7, 8]. Additionally, a minor tolerance threshold $\epsilon_{\text{quant}}$ could allow for small fluctuations due to rounding.
6 Empirical Cost Estimates: GPT-3 and Gemma 3
6.1 Model Sizes in Bytes
The storage requirements for different model sizes vary depending on the number of parameters and the level of quantization. The table below summarizes the size estimates for two models—GPT-3 and Gemma 3—at 32-bit full precision and 4-bit quantization:
Model 32-bit (Full Precision) 4-bit Quantized
GPT-3 (175B parameters) $\approx 700$ GB $\approx 87.5$ GB
Gemma 3 (27B parameters) $108$ GB $13.5$ GB
Table 1: Model sizes for 32-bit and 4-bit quantization schemes.
Hence, the 4-bit version is significantly more compact.
6.2 Ratio of Computation: Miners vs. Verifiers
Formally, let $C_{\text{train}}$ denote the computational cost (in GPU-hours or a similar metric) for a miner to perform one gradient-descent update on a given mini-batch. This cost typically includes the forward pass, backward pass, and weight-update step in full 32-bit floating-point arithmetic:
$C_{\text{train}}\;=\;C_{\text{forward}}^{(32)}+C_{\text{backward}}^{(32)}+C_{\text{update}}^{(32)}.$
By contrast, verifiers in PoGO only need to check two main components of the proposed block:
  1. 1.
    Quantized Loss Check (Forward Pass).
    A verifier must confirm that the 4-bit quantized model $\widetilde{\theta}_{t+1}$ indeed achieves the claimed lower loss on a small subset of the dataset (drawn from a public VRF seed). Denote this verification subset as $\mathcal{D}_{\text{ver}}\subset\mathcal{D}$, where $\lvert\mathcal{D}_{\text{ver}}\rvert\ll\lvert\mathcal{D}\rvert$.
    The cost of this forward pass can be written as
    $C_{\text{forward}}^{(4)}(\mathcal{D}_{\text{ver}}),$
    where the superscript $(4)$ indicates 4-bit precision. Empirically, modern GPU tensor cores can process 4-bit representations up to 8×8$\times$8 × faster than 32-bit [7, 9]. If $\alpha=\frac{\lvert\mathcal{D}_{\text{ver}}\rvert}{\lvert\mathcal{D}\rvert}$ is the fraction of the entire training data used for verification, the verifier’s forward-pass cost is approximately
    $C_{\text{forward}}^{(4)}(\alpha)\;\approx\;\alpha\;\frac{C_{\text{forward}}^{(32)}(\mathcal{D})}{8},$
    assuming the time scales linearly with the dataset size and an 8×8$\times$8 × speedup going from 32-bit to 4-bit.
  2. 2.
    Merkle Leaf Check.
    The verifier must also check a random leaf from the Merkle tree of the miner’s proposed 32-bit model. Let $C_{\text{merk}}$ denote the cost of the following operations:
    1. (a)
      retrieving a leaf value $\texttt{leaf}_{i}$;
    2. (b)
      verifying the Merkle path;
    3. (c)
      comparing the retrieved value against the quantized version.
    Since only a single (or a few) leaves are sampled, $C_{\text{merk}}$ is effectively constant per block—independent of model size $d$ or dataset size. In practice, this cost is dominated by lightweight hash operations and a simple numerical comparison between 4-bit and 32-bit weights (allowing for rounding).
Hence, the total verification cost per block, $C_{\text{verify}}$, can be modeled as
$C_{\text{verify}}\;=\;C_{\text{forward}}^{(4)}(\alpha)\;+\;C_{\text{merk}}.$
In contrast, the miner’s cost $C_{\text{train}}$ involves the more expensive 32-bit forward and backward passes over the full mini-batch, plus other overheads (e.g., optimizer steps):
$C_{\text{train}}\;\approx\;C_{\text{forward}}^{(32)}(\mathcal{D})\;+\;C_{\text{backward}}^{(32)}(\mathcal{D})\;+\;C_{\text{update}}^{(32)}.$
Because $\alpha\ll 1$ and 4-bit computations can be significantly faster than 32-bit, we typically have
$C_{\text{verify}}\;\ll\;C_{\text{train}}.$
For instance, even if $C_{\text{forward}}^{(32)}(\mathcal{D})$ accounts for 10 GPU-hours in training, using $\alpha=0.01$ (just 1% of the data for verification) and an 8×8$\times$8 × speedup yields
[…]
plus a negligible $C_{\text{merk}}$ for the Merkle leaf check. Thus, $C_{\text{verify}}$ might easily be 10similar-toabsent10$\sim$ 10∼ 10100×100$\times$100 × cheaper than $C_{\text{train}}$, depending on the dataset fraction $\alpha$ and hardware efficiency. For very large models (billions of parameters), this gap can grow further, since most verification overhead (the leaf-check) remains essentially constant, while the training cost scales with the model and batch size.
In short, PoGO’s design ensures that the heaviest computational burden lies on the miners who must perform genuine training. Meanwhile, verifiers perform relatively lightweight checks on a small subset of data in 4-bit precision and confirm Merkle proofs for random leaves, thereby maintaining high security at a much lower cost.
7 Longer Block Times
Unlike standard blockchains with block times measured in seconds or minutes, PoGO might push block times to hours (or even a day) to accommodate meaningful training steps. This ensures:
  • Miners have sufficient time to do non-trivial gradient descent on large mini-batches.
  • The network participants can download or partially download the 4-bit model from IPFS within the finalization window $w$ (e.g., 20 blocks, each possibly an hour).
Although this is a departure from fast finality blockchains, it may be acceptable in specialized ML-training blockchains where throughput is less critical than the correctness and authenticity of the training steps.
8 Binary Diffs and Incremental Updates
An alternative to publishing the entire model after each update is to only publish binary diffs (the difference between consecutive 4-bit versions). This can reduce the size of each update, but for large deep learning models, the fraction of weights that change significantly in each step can still be large. Empirically, while binary diffs can yield some savings, it is often not a dramatic order-of-magnitude improvement [7]. Hence, it is a small optimization that can be optionally employed to reduce bandwidth and storage overhead.
9 Fine-Tuning Use Case
The same PoGO mechanism applies for fine-tuning a pretrained model. The only difference is that the dataset is now user-provided or domain-specific. Verifiers still:
  1. 1.
    Train with mini-batch provided by the model owner or paying user (possibly the entire user dataset if small).
  2. 2.
    Check that the new 4-bit model lowers the loss:
    ^$\widehat{\mathcal{L}}(\widetilde{\theta}_{\text{fine}+1})\;<\;\widehat{\mathcal{L}}(\widetilde{\theta}_{\text{fine}})-\epsilon_{\text{fine}},$
    on that mini-batch or a separate held-out subset.
  3. 3.
    Then request a random leaf from the new full 32-bit parameters to confirm consistency via Merkle proofs.
Since fine-tuning can overfit if the dataset is small, the random mini-batch might be the same data used for training. As long as the protocol is transparent about which data is used and the improvement is validated, it remains consistent with PoGO principles.
10 Security Analysis
10.1 Two-Phase Randomness and Merkle Root
Publishing the Merkle root at block $N$ but only revealing a random leaf at block $N+(w/2)$ prevents the miner from backdating or selectively generating a forged Merkle tree or quantized model. Because the seed for random leaf selection is only known later, the miner cannot guess which leaf to prepare with spurious data.
10.2 Data Availability and Attestations
Large data (4-bit or 32-bit) is stored off-chain (e.g., IPFS). The network relies on:
  1. 1.
    Partial checks (4-bit model hash, random leaf check).
  2. 2.
    Attestations: verifiers who can access and verify the data sign a positive attestation if it is consistent, or a negative attestation if they detect a mismatch (e.g., the Merkle leaf is invalid, the quantized model does not truly reduce loss, or data was not provided) [4].
If not enough positive attestations appear by block $N+w$ (e.g., fewer than $2/3$ of stake), or if many negative attestations are submitted, the update is rejected and the miner is slashed.
11 Incentives and Slashing
  • Rewards: Each confirmed update yields a training reward (paid by the model owner or from protocol incentives) that goes mostly to the single random miner of that block, with a fraction (proportional to their stake) to verifiers who produce timely attestations.
  • Slashing: If a miner fails to publish valid data or if sufficient negative attestations show an inconsistency, the block $N$ update is rejected at block $N+w$ and the miner’s stake is slashed.
This encourages honest participation. Because all miners who are not the leader for block $N$ become verifiers, the system harnesses the entire staking base for robust checks. Negative attestations serve as direct evidence of misconduct, allowing the final aggregator to impose penalties.
12 The Role of Staking and Alternative Competitive Designs
Staking plays a critical role in this system beyond just slashing and rewards. We rely heavily on staking to enable randomness in selecting both the leader miner for each block and the next machine learning model to be trained. This randomness is essential to prevent manipulation and ensure fair participation.
One might imagine an alternative design that does not rely on staking at all. In such a system, there could be a single global model [10], and any participant can attempt to train it. Miners would compete openly, and the winner would be the one who most successfully lowers the model’s loss function,
$\mathcal{L}(\theta),$
on a fixed validation set. A block that claims a model $\theta^{*}$ is accepted only if
$\mathcal{L}(\theta^{*})\;<\;\mathcal{L}(\theta_{\text{previous}})-\epsilon,$
for some threshold $\epsilon$. The “best” fork is then chosen as the one achieving the lowest loss so far.
While this approach eliminates the need for randomness/staking, it also forces the protocol to focus on a single model architecture and task at a time. Our staked PoGO design generalizes to multiple concurrent models, ensures random selection of tasks, and provides a robust mechanism for slashing dishonest updates.
13 Comparison to Bittensor Approaches
Bittensor v1: A Peer-to-Peer Intelligence Market.
The original Bittensor framework [2] proposed a decentralized “intelligence market,” wherein peers (each hosting a neural model) directly evaluate one another’s outputs and assign weights that reflect perceived utility or “information-theoretic” value. Nodes receive additional stake if they are deemed valuable by others, creating an incentive to provide useful model outputs. This peer-to-peer approach is elegant in that it relies on local pairwise evaluations rather than a centralized oracle. However, as the Bittensor team notes, it also introduces a risk of collusion or sybil-like attacks if adversarial subgroups artificially inflate one another’s scores. Their mechanism partially mitigates this by rewarding only weights recognized by a majority share of the network, though in practice it remains challenging to definitively prove that a model’s output is genuinely useful beyond local evaluations.
Bittensor v2: Stake-Based Consensus for Utility Scoring.
A subsequent version of Bittensor [2] delves deeper into an incentive function that combines network “consensus” checks with pairwise utility signals. This version formalizes a two-team (honest vs. cabal) game to analyze how stake evolves under different weight-assignment strategies. The design aims to penalize nodes that assign obviously skewed weights by limiting their overall stake growth if a majority of honest nodes disagree. While this improves collusion resistance compared to simple local weighting, the protocol’s complexity increases, and it still relies on collectively agreeing that certain weights are “correct” or “incorrect” in an inherently subjective, model-to-model sense.
Critique and Comparison.
Both Bittensor versions share the vision of transforming costly computations (in their case, model inference or representational learning) into a decentralized marketplace of machine intelligence. However, PoGO differs in scope and methodology by providing a cryptographically verifiable proof of actual training progress, rather than primarily relying on peer-based agreement about outputs’ quality. In Bittensor, a node’s “usefulness” is established by neighbors’ feedback signals, which can be manipulated if dishonest peers coordinate. Meanwhile, PoGO enforces a more direct measure of work: each block must present evidence that it has lowered a publicly verifiable loss on a known dataset.
We view these lines of research as complementary rather than directly competing. Bittensor’s approach excels at building an ecosystem of diverse models that can discover and reward novel capabilities, whereas PoGO is well-suited for verifying large-scale training (e.g. fine-tuning or full model updates) using quantized gradients and Merkle proofs. Future hybrid systems might incorporate Bittensor-style local scoring within PoGO’s cryptographic framework to more richly reward high-performing sub-networks while still preserving strong security against collusion.
14 Additional Incentive Mechanisms:
Storage Rental and Dynamic Price Governance
While Sections 9–10 describe basic rewards and slashing, PoGO also incorporates a storage rental mechanism to ensure that models do not linger cost-free on the network. In particular, uploading a new model or fine-tuning dataset requires users to commit a certain amount of POGO tokens to rent storage for some number of blocks. If this rental expires, nodes need not store or attest to that model.
14.1 Renting Storage for Models
When a user issues an UploadModel transaction, they:
  • Provide a hash of the model (and possibly partial data).
  • Commit POGO tokens to cover rentedBlocks of storage.
  • Within a modelUploadWindow, they must upload the model to IPFS or a similar network. If they fail to do so, verifiers can attest that the model is unavailable, and the system rejects it.
A user can top up storage later (e.g. via TopupStorageRental) if training or forking continues beyond the initially funded period. If no top-up is provided by the time storage expires, other nodes may discard the model data.
14.2 Consensus-Driven Price Adjustments
PoGO maintains a dynamic per-block price for storage (per GB per block) and for compute (per training step). We call these:
[…]
Each consensus leader can nudge these prices by a small ±$\pm\Delta$ (e.g. up to 0.01% per block), capturing supply-demand fluctuations. Thus, if network usage surges, leaders will collectively trend the price upward; if usage declines, it trends downward. This dynamic governance ensures a rough market equilibrium without external oracles.
14.3 Forking and Fine-Tuning Fees
Any user can fork an existing model (ForkModel transaction) by paying a new round of storage rental. Fine-tuning tasks also require committing POGO tokens to pay for compute usage, with a separate fraction of basicComputePrice possibly discounted (e.g., a fineTuningFraction parameter if the dataset is smaller).
14.4 Incentive Impact
This approach aligns each participant’s economic incentives:
  • Model Uploaders pay for both storage time and future training steps; if they fail to provide data, they lose their deposits.
  • Miners (Block Leaders) earn block rewards plus fees from performing the gradient optimization, scaled by basicComputePrice.
  • Consensus (Stakeholders) collaboratively adjusts storage/compute prices in small increments, ensuring the network remains neither overburdened nor underpriced.
Overall, these dynamics encourage stable, long-term use of PoGO for large-scale training or fine-tuning tasks, with deferred publication and rental top-ups ensuring data availability only so long as it is economically justified.
15 Conclusion and Future Directions
We have presented PoGO, enhancing earlier Proof of Gradient Optimization with several key ideas. These include quantized models (4-bit) [7, 8] to drastically reduce bandwidth and compute overhead for verification, and a Merkle proof of the full-precision model to ensure no hidden tampering with high-precision parameters. We introduced positive and negative attestations, enabling quick rejection (and slashing) if verifiers detect dishonest updates. Our empirical cost estimates demonstrate that verifying large models—such as GPT-3 or Gemma 3—is viable, with verification costs an order of magnitude lower than training. We also acknowledged extended block times as a design trade-off to accommodate meaningful ML training, and ensured fine-tuning compatibility with minimal changes to the verification pipeline.
Future work includes exploring advanced zero-knowledge proofs for partial gradient verification and refining quantization strategies for even greater efficiency. PoGO opens up new horizons where blockchain security and real-world ML training can reinforce each other, providing a decentralized ecosystem for building, verifying, and sharing large models.
Acknowledgments
We thank the broader blockchain and AI communities for early discussions, and the reviewers of earlier PoGO drafts for foundational concepts.